Your local athletic center is planning a $500,000 expansion to its current facility. This cost will be depreciated on a straight-line basis over a 20-year period. The expanded area is expected to generate $175,000 in additional annual sales. Variable costs are 32 percent of sales, the annual fixed costs are $40,000, and the tax rate is 21 percent. What is the operating cash flow for the first year of this project?
Answer #1

Annual depreciation = 500,000 / 20 = 25,000

variable costs = 175,000 * 0.32 = 56,000

Operating cash flow = (Sales - variable costs - fixed costs - depreciation)(1 - tax) + depreciation

Operating cash flow = (175,000 - 56,000 - 40,000 - 25,000)(1 - 0.21) + 25,000

Operating cash flow = 42,660 + 25,000

Operating cash flow = $67,660